Using Intermediate Value Theorem, show that


x2 = √𝑥 + 1 has a root in (1,2)


1
Expert's answer
2022-03-19T02:39:36-0400

Solution


Given that


"x^2=\sqrt{x+1}"


This can be re-written as


"f(x)=x^2-\sqrt{x+1}" where "f(x)=0" will give us the root(s) of the equation.



Now for


"x=1" "f(1)=(1)^2-\sqrt{(1)+1}=-0.4142... <0"


And


"x=2" "f(2)=(2)^2-\sqrt{(2)+1}=2.2679... >0"


Since "f(1)<0" and "f(2)>0"


Therefore, "f(x)=0" for some value of "x" which lies with in the interval "(1, 2)"


Hence "x^2=\sqrt{x+1}" has a root within the interval "(1, 2)"


The plot below shows that "x^2=\sqrt{x+1}" has root between "(1, 2)" , which is "x=1.221"